# Quantum property testing in sparse directed graphs

### v2

@@ 2.1 Notations and basic definitions
Let us denote $[n]=\{1,\dots,n\}$ and $[n]_{0}=\{0,\dots,n\}$. When dealing with Boolean variables, we will usually use $b\in\{-1,1\}$ instead of $b\in\{0,1\}$. We can get to one from the other easily with the mapping $b\mapsto 1-2b$, or its inverse, which means that $-1$ is going to be treated as the “true” or “accepting” value. The reason for using $\{-1,1\}$ is that when dealing with dual polynomials, it is easier to use this notation.
We denote by $1^{n}$ the length-$n$ binary vector made only of $1$s, and respectively $-1^{n}$. The Hamming weight $|x|$ of $x\in\{-1,1\}^{n}$ is then defined as the number of $-1$s in $x$, that is $|x|=\#\{i\in[n]:x_{i}=-1\}$. Let $H_{\leq w}^{n}=\{x\in\{-1,1\}^{n}:|x|\leq w\}$ denote the set of length-$n$ binary vectors with Hamming weight at most $w$. For any $x\in\mathbb{R}$, $\mathrm{sgn}(x)=1$ when $x\geq 0$, and $-1$ otherwise.
@@
Finally, throughout this work, notations $O(\cdot)$ and $\Omega(\cdot)$ will be hiding the dependencies on parameters $\varepsilon$, $k$ and $d$ that we consider to be constants. Additionally, we will use $\widetilde{O}(\cdot)$ and $\widetilde{\Omega}(\cdot)$, where we hide poly-logarithmic dependencies on the parameters.
@@ 2.2 Query complexity
$\mathcal{O}_{G}(v,i)=\begin{cases}w,&\text{if $w\in V$ is the $i$-th neighbor of $v$};\\ \bot,&\text{if deg}(v)<i.\end{cases}$
@@ 2.3 Property testing
Let $10<\varepsilon<$ 1 be a constant. An algorithm $\mathcal{A}$ is an $\varepsilon$-tester for the property $\mathcal{P}$ if
@@ 2.4 Grover search
We are not going into details about quantum computing, because for this paper, it suffices to state very few results in the field and use them in a black box way. One of the most important results in quantum computing is Grover’s algorithm for unordered search [Grover]: finding a marked element in an unordered database of size $N$ takes $\Theta(N)$ queries classically, but the quantum query complexity of the same task is $\Theta(\sqrt{N})$. We are going to use a particular variant of this result that has been used many times in the literature (see e.g., [Ambainis_search, Item 3 in Section 2.2], which was implicitly proved in [BBHT_search]).
@@
Let $1\leq t_{0}\leq N$. There exists a quantum algorithm that, given $t_{0}$ and query access to any function $f:S\to\{0,1\}$, makes $O(\sqrt{N/t_{0}})$ queries to $f$ and outputs either “not found” or an element uniformly at random in $f^{-1}(1)$. Moreover, when $|f^{-1}(1)|\geq t_{0}$, the latter occurs with high constant probability.
@@ 2.5 Problem definitions
As discussed in the introduction (Section 1.3), very little was known about the property testing version of this problem prior to this work. We only know the that the complexity is $\Theta(N^{1/3})$ when $k=2$, and it is between $\Omega(N^{1/3})$ and $O\left(N^{\frac{1}{2}\left(1-\frac{1}{2^{k}-1}\right)}\right)$ for larger $k$.
@@
The problem of $\varepsilon$-testing $k$-collision-freeness of a sequence from $[R]^{N}$ can be reduced to $\frac{\varepsilon N}{d(N+R)}$-testing $k$-star-freeness of an $(N+R)$-vertex sparse directed graph with out-degree bound $d\geq 1$.
@@ 3 Quantum algorithm for testing subgraph-freeness
We will use Breadth-first search (BFS) in order to explore $G$ layer by layer: starting from a given vertex, first it explores its direct (out-)neighbors, then their unexplored (out-)neighbors etc. We will run BFS up to some limited depth $\ell$, so that the $-\ell$ BFS algorithm has query complexity at most $d^{\ell}$ where $d$ is the maximum (out-)degree of $G$. In our application, $d$ and $\ell$ are constants, and thus the query complexity is $d^{\ell}=O(1)$.
@@ 3.1 The algorithm for $k=2$
[…]
@@
[…]
@@ 3.2 The algorithm for general $k$
Theorem 3.3 (Restatement of Theorem 1.1).
@@
In order to extend the $k=2$ case described above to larger $k$, we first try to find many partial $H$-instances with $k-1$ source components found, and then extend one of them to a complete $H$-instance. We present a brief description of our algorithm below, where $h$ is the number of vertices of $H$:
@@
[…]
@@
      Perform a Grover search $t_{i}$ times on the vertices $V\setminus\mathcal{S}_{i-1}^{\prime}$ in the following way. A vertex $v$ is marked if there exist $i-1$ other vertices $u_{j}\in\mathcal{S}_{j}$ for each $j\in[i-1]$ such that $u_{1},\dots,u_{i-1}$ and $v$ are from $i$ different source components of an $H$ subgraph of $G$. If we do not find $t_{i}$ vertices like this, output Reject, otherwise let $\mathcal{S}_{i}$ denote the set of the vertices $v$ that we found.
@@
    Perform Grover search on $V\backslash\mathcal{S}_{k-1}^{\prime}$ to find a complete $H$-instance. I.e., a vertex $v$ is marked if there exist $k-1$ other vertices $u_{j}\in\mathcal{S}_{j}$ for each $j\in[k-1]$ such that $u_{1},\dots,u_{k-1}$ and $v$ are from the $k$ different source components of an $H$ subgraph of $G$.
@@
The correctness proof is similar to the $k=2$ case. Proposition 3.2 tells us that in $\mathcal{S}_{1}$ there are $\Omega(t_{1})$ many vertices that are from a source component of an $H$-copy. Because of the source-disjointness of the $H$-copies, when $i=2$, there are $\Omega(t_{1})$ many 1-partial solutions that can be extended to a complete $H$ instance by disjoint remaining source components. As Grover search provides uniformly random marked elements, a constant fraction of the $t_{2}$ many 2-partial solutions are actually extendable to $H$ in a similar, disjoint way. This continues to be true in each iteration: (with high probability) a constant fraction of the $t_{i-1}$ many $(i-1)$-partial solutions are extendable to complete $H$ instances by disjoint remaining source components. This way, the last step is going to find an $H$-subgraph with high probability.
[…]
@@ 4 Collision-freeness lower bound
As discussed in Section 2, we are going to prove a lower bound on the problem of testing $k$-collision-freeness.
Theorem 4.1 (Restatement of Theorem 1.3).
@@
The proof of the theorem is at the end of Section 4.3. Observe that Theorem 1.2 is implied by Theorem 4.1 and the reduction in Proposition 2.6. Our proof mostly follows the structure of [polynomial_strikes_back, Section 6.1], and in particular it uses the notion of dual polynomial for non-Boolean partial symmetric functions. Our main technical contribution in this section is the proof of Lemma 4.18, because the corresponding proof in [polynomial_strikes_back] crucially relies on a fact that does not hold for our problem. We will discuss it in detail below.
@@ 4.1 The (dual) polynomial method
We now consider a property of a sequence of integers as a function $F:D\subseteq[R]_{0}^{N}\to\{-1,1\}$. The symbol 00 will play a special role that will be exhibited later on. Unfortunately one cannot just take the polynomial of those integers. The standard approach (see [Aaronson02]) is to encode $s=(s_{1},\dots,s_{N})\in[R]_{0}^{N}$ into binary variables $x=(x_{i,j})_{i\in[N],j\in[R]_{0}}\in\{-1,1\}^{N(R+1)}$ encoding whether $s_{i}=j$ as follows: $x_{i,j}=-1$ if $s_{i}=j$, and $x_{i,j}=1$ otherwise. Let $H_{b}^{N(R+1)}\subseteq\{-1,1\}^{N(R+1)}$ be the set of all possible encodings of vectors $s$, that is, for every $i\in[N]$ there is exactly one $j\in[R]_{0}$ such that $x_{i,j}=-1$.
@@
But before doing this, we are going to relax the constraints on the domain $D_{b}$ in the case of symmetric functions, while we decrease its dimension. When $F$ is symmetric (i.e., $F(s)=F(s\circ\pi_{N})$ for any permutation $\pi_{N}$ of $[N]$), one can instead define a function $F_{\leq N}$ with weaker constraints by removing the variables corresponding to the symbol 00. Define $H_{\leq N}^{NR}$ as the set of length-$(NR)$ binary vectors with Hamming weight at most $N$. Given any $x\in H_{\leq N}^{NR}$, we define its frequency vector $z(x)=(z_{0},z_{1},\dots,z_{R})$ with $z_{j}=\#\{i:x_{ij}=-1\}$, for $1\leq j\leq R$, and $z_{0}=N-z_{1}-\ldots-z_{R}$. From the vector $z(x)$, one can define a valid sequence of integers $s(x)\in[R]_{0}^{N}$: it can be any sequence from $[R]_{0}^{N}$ that has frequency vector $z(x)$. Now we can define $F_{\leq N}$ on domain $D_{\leq N}$ as
@@
Lemma 4.7 ([Aaronson02, ambainis_poly],[BT20, Theorem 3.9]).
@@
Let $F:D\subseteq[R]_{0}^{N}\to\{-1,1\}$ be symmetric. Define $F_{\leq N}:D_{\leq N}\to\{-1,1\}$ as above. Then ${\mathrm{dpdeg}}_{\delta}(F_{\leq N})\geq\Delta$ iff there exists a function $\psi:\{-1,1\}^{NR}\to\mathbb{R}$ such that
@@ 4.2 Preparation
Technically, the problem we use in the proof of Theorem 4.1 is slightly more restricted than $k$-collision-freeness: we want to distinguish no $k$-collision from many distinct collisions of size at least $k$.
@@
Let $\gamma\in(0,1)$. The symmetric function $\mathrm{Collision}^{k,\gamma}_{N,R}:D_{\mathrm{Collision}^{k,\gamma}_{N,R}}\subset[R]^{N}\to\{-1,1\}$ is defined by $\mathrm{Collision}^{k,\gamma}_{N,R}(s)=-1$ if no integer occurs at least $k$ times in $s$, $\mathrm{Collision}^{k,\gamma}_{N,R}(s)=1$ if there are more than $\gamma R$ distinct integers that occur at least $k$ times in $s$, and it is undefined otherwise.
@@
To prove a bound on the $\mathrm{Collision}$ function, we will actually relate it to the composition of two more elementary functions $g,h$, where by composition we mean $(g\circ h)(x)=g(h(x_{1}),\dots,h(x_{n}))$ (where $x=(x_{1},\dots,x_{n})$ and each $x_{i}$ is a binary vector of appropriate dimension). Let us define (i) the threshold function $\mathrm{THR}_{N}^{k}:\{-1,1\}^{N}\to\{-1,1\}$ which is $-1$ if the input bitstring contains at least $k$ many $-1$s, and it is 1 otherwise; and (2) the gap version of $\mathrm{OR}$, that is $\mathrm{GapOR}_{R}^{\gamma}:D_{\mathrm{GapOR}_{R}^{\gamma}}\subset\{-1,1\}^{R}\to\{-1,1\}$ which takes value 1 if the input is $1^{R}$, $-1$ if the input contains at least $\gamma R$ many $-1$s, and is undefined otherwise. We show that the double-promise approximate degree of $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ lower bounds the quantum query complexity of the collision problem.
@@
Let $k\geq 3$, $10<\gamma<$ 1, $\delta>0$ and $c>2$ be constants such that $N/c\leq R\leq N/2$. If the double-promise $\delta$-approximate degree of $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ on domain further restricted to $H_{\leq N}^{NR}$ is at least $\Delta$, then every quantum algorithm computing $\mathrm{Collision}^{k,\gamma/c}_{N,N}$ with error $\delta/2$ must require at least $\Delta/2$ queries.
Before proving this lemma, we prove some helper propositions. In order to apply the dual polynomial method for partial symmetric functions, we start by proving that $\mathrm{Collision}^{k,\gamma^{\prime}}_{N,R^{\prime}}$ is at least as hard as a very similar problem. We introduce a “dummy-augmented” version $\mathrm{dCollision}_{N,R}^{k,\gamma}:D_{\mathrm{dCollision}_{N,R}^{k,\gamma}}\subseteq[R]_{0}^{N}\to\{-1,1\}$ of the problem $\mathrm{Collision}_{N,R}^{k,\gamma}$ for the purpose of proving Lemma 4.10, where now the input sequence can have integer 00, but those 00s are just ignored when they occur. We show that it is enough to prove a lower bound for this second version.
@@
Let $k\geq 3$, $10<\gamma<$ 1 and $c>2$ be constants such that $N/c\leq R\leq N/2$. Then $\mathrm{dCollision}^{k,\gamma}_{N,R}$ can be reduced to $\mathrm{Collision}^{k,\gamma/c}_{N,N}$.
@@
An input to $\mathrm{dCollision}^{k,\gamma}_{N,R}$ is a sequence $s=(s_{1},\dots,s_{N})$ where each $s_{i}\in[R]_{0}$. Let us define a family of functions $T_{i}$ that map from $[R]_{0}$ to $[R^{\prime}]$ for $R^{\prime}=R+\lceil N/2\rceil$: $T_{i}(s)=s$ if $s>0$ and $T_{i}(0)=R+\lceil i/2\rceil$.
@@
We are now ready to give the proof of Lemma 4.10.
Proof of Lemma 4.10.
By Proposition 4.11, instead of $\mathrm{Collision}^{k,\gamma/c}_{N,N}$ we can consider $\mathrm{dCollision}^{k,\gamma}_{N,R}$ (with the appropriate parameters) to show a lower bound. By Proposition 4.12, to relate the query complexity of $\mathrm{dCollision}^{k,\gamma}_{N,R}$ to the double-promise degree of $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ with domain further restricted to $H_{\leq N}^{NR}$. ∎

### v3

@@ 2.1 Notations and basic definitions
Let us denote $n]=\{1,\dots,n\}$ and $n]_{0}=\{0,\dots,n\}$. When dealing with Boolean variables, we will usually use $b\in\{-1,1\}$ instead of $b\in\{0,1\}$. We can get to one from the other easily with the mapping $b\mapsto 1-2b$, or its inverse, which means that 1-1 is going to be treated as the “true” or “accepting” value. The reason for using $\{-1,1\}$ is that when dealing with dual polynomials, it is easier to use this notation.
We denote by $1^{n}$ the length-$n$ binary vector made only of $1$s, and respectively $1^{n}$. The Hamming weight $|x|$ of $x\in\{-1,1\}^{n}$ is then defined as the number of 1-1s in $x$, that is $|x|=\#\{i\in[n]:x_{i}=-1\}$. Let $H_{\leq w}^{n}=\{x\in\{-1,1\}^{n}:|x|\leq w\}$ denote the set of length-$n$ binary vectors with Hamming weight at most $w$. For any $x\in\mathbb{R}$, $\mathrm{sgn}(x)=1$ when $x\geq 0$, and 1-1 otherwise.
@@
[…]
@@ 2.2 Query complexity
[…]
@@ 2.3 Property testing
Let $0<\varepsilon<1$ be a constant. An algorithm $\mathcal{A}$ is an $\varepsilon$-tester for the property $\mathcal{P}$ if
@@ 2.4 Grover search
We are not going into details about quantum computing, because for this paper, it suffices to state very few results in the field and use them in a black box way. One of the most important results in quantum computing is Grover’s algorithm for unordered search [Grover]: finding a marked element in an unordered database of size $N$ takes $\Theta(N)$ queries classically, but the quantum query complexity of the same task is $\Theta(\sqrt{N})$. We are going to use a particular variant of this result that has been used many times in the literature (see e.g., [Ambainis_search, Item 3 in Section 2.2], which was implicitly proved in [tight_lb_grover]).
@@
Let $1\leq t_{0}\leq N$. There exists a quantum algorithm that, given $t_{0}$ and query access to any function $f:S\to\{0,1\}$, makes $O(\sqrt{N/t_{0}})$ queries to $f$ and outputs either “not found” or an element uniformly at random in $f^{-1}(1)$. Moreover, when $|f^{-1}(1)|\geq t_{0}$ the later occurs with high constant probability.
@@ 2.5 Problem definitions
As discussed in the introduction (Section˜1.3), very little was known about the property testing version of this problem prior to this work. We only know the that the complexity is $\Theta(N^{1/3})$ when $k=2$, and it is between $\Omega(N^{1/3})$ and $O\left(N^{\frac{1}{2}\left(1-\frac{1}{2^{k}-1}\right)}\right)$ for larger $k$.
@@
The problem of $\varepsilon$-testing $k$-collision-freeness of a sequence from $R]^{N}$ can be reduced to $\frac{\varepsilon N}{d(N+R)}$-testing $k$-star-freeness of an $(N+R)$-vertex sparse directed graph with out-degree bound $d\geq 1$.
@@ 3 Quantum algorithm for testing subgraph-freeness
We will use Breadth-first search (BFS) in order to explore $G$ layer by layer: starting from a given vertex, first it explores its direct (out-)neighbors, then their unexplored (out-)neighbors etc. We will run BFS up to some limited depth $\ell$, so that the depth-$\ell$ BFS algorithm has query complexity at most $d^{\ell}$ where $d$ is the maximum (out-)degree of $G$. In our application, $d$ and $\ell$ are constants, and thus the query complexity is $d^{\ell}=O(1)$.
@@ 3.1 The algorithm for $k=2$
    Sample a uniformly random vertex subset $\mathcal{S}$ of size $t=\Theta(N^{1/3})$ in $G$. Perform a depth-$h$ BFS from every vertex in $\mathcal{S}$.
@@
By Proposition˜3.2, with high probability, a constant fraction of the $t$ vertices in $\mathcal{S}$ are part of a source component in source-disjoint $H$-subgraphs of $G$. For such vertices, the BFS in step 1 will discover the entire source component, as well as all other vertices reachable from that source component in $H$. Then, in step 2 we search for a vertex that is in the remaining source component of such an instance of $H$ that we already partly discovered. This can be verified by doing a depth-$h$ BFS from it and checking if this completes an $H$-instance with one of the previously sampled vertices’ neighborhoods. As we mentioned, by Proposition˜3.2 with high probability there are $\Omega(t)$ many marked vertices. This proves the correctness of the algorithm.
Finally, we bound the algorithm’s query complexity. Step 1 makes $O(t)=O(N^{1/3})$ many (classical) queries. In step 2 we use Theorem˜2.2 and Remark˜2.3: checking whether a vertex is marked requires running a depth-$h$ BFS from it, which costs $c=O(1)$ queries. We argued that there are $\Omega(t)$ many marked vertices, so Grover search makes $O(\sqrt{N/t})=O(N^{1/3})$ quantum queries.
@@ 3.2 The algorithm for general $k$
Theorem 3.3 (Restatement of Theorem˜1.1).
@@
In order to extend the $k=2$ case described above to larger $k$, we first try to find many partial $H$-instances with k1k-1 source components found, and then extend one of them to a complete $H$-instance. We present a brief description of our algorithm below, where $h$ is the number of vertices of $H$:
@@
    Sample a uniformly random vertex subset $\mathcal{S}_{1}$ in $G$ of size $t_{1}$. Perform a depth-$h$ BFS from every vertex in $\mathcal{S}_{1}$. Let $\mathcal{S}_{1}^{\prime}=\mathcal{S}_{1}$.
@@
      Perform a Grover search $t_{i}$ times on the vertices $V\setminus\mathcal{S}_{i-1}^{\prime}$ in the following way. A vertex $v$ is marked if there exist i1i-1 other vertices $u_{j}\in\mathcal{S}_{j}$ for each $j\in[i-1]$ such that $u_{1},\dots,u_{i-1}$ and $v$ are from $i$ different source components of an $H$ subgraph of $G$. If we do not find $t_{i}$ vertices like this, output Reject, otherwise let $\mathcal{S}_{i}$ denote the set of the vertices $v$ that we found.
@@
    Perform Grover search on $V\backslash\mathcal{S}_{k-1}^{\prime}$ to find a complete $H$-instance. I.e., a vertex $v$ is marked if there exist k1k-1 other vertices $u_{j}\in\mathcal{S}_{j}$ for each $j\in[k-1]$ such that $u_{1},\dots,u_{k-1}$ and $v$ are from the $k$ different source components of an $H$ subgraph of $G$.
@@
The correctness proof is similar to the $k=2$ case. Proposition˜3.2 tells us that in $\mathcal{S}_{1}$ there are $\Omega(t_{1})$ many vertices that are from a source component of an $H$-copy. Because of the source-disjointness of the $H$-copies, when $i=2$, there are $\Omega(t_{1})$ many 1-partial solutions that can be extended to a complete $H$ instance by disjoint remaining source components. As Grover search provides uniformly random marked elements, a constant fraction of the $t_{2}$ many 2-partial solutions are actually extendable to $H$ in a similar, disjoint way. This continues to be true in each iteration: (with high probability) a constant fraction of the $t_{i-1}$ many (i1)(i-1)-partial solutions are extendable to complete $H$ instances by disjoint remaining source components. This way, the last step is going to find an $H$-subgraph with high probability.
To bound the query complexity, first note that in every application of Grover search, checking whether a vertex is marked (depth-$h$ BFS) takes $O(1)$ queries. The first iteration’s Grover searches find $t_{2}$ partial $H$-instances with $2$ of its source components found, which takes $O(t_{2}\sqrt{N/t_{1}})$ queries (by Theorem˜2.2). Similarly, for $i$-th iteration there are $\Omega(t_{i-1})$ marked elements (see the previous argument), so the algorithm performs $O(t_{i}\sqrt{N/t_{i-1}})$ quantum queries for every $i\in[k-1]$. Finally, finding one complete $H$-instance costs $O(\sqrt{N/t_{k-1}})$ queries. Thus the total query complexity is $O(t_{1}+\sum_{i=1}^{k-1}t_{i+1}\sqrt{N/t_{i}})$ with $t_{k}=1$. Similar to the multi-collision algorithm in [zhandry_collision, Section 3], we can equate all terms by setting $t_{i}=\Theta\Big(N^{\frac{2^{k-i}-1}{2^{k}-1}}\Big)$, which yields the final quantum query complexity $O\Big(N^{\frac{1}{2}\left(1-\frac{1}{2^{k}-1}\right)}\Big)$.
@@ 4 Collision-freeness lower bound
As discussed in Section˜2, we are going to prove a lower bound on the problem of testing $k$-collision-freeness.
Theorem 4.1 (Restatement of Theorem˜1.3).
@@
The proof of the theorem is at the end of Section˜4.3. Observe that Theorem˜1.2 is implied by Theorem˜4.1 and the reduction in Proposition˜2.6. Our proof mostly follows the structure of [polynomial_strikes_back, Section 6.1], and in particular it uses the notion of dual polynomial for non-Boolean partial symmetric functions. Our main technical contribution in this section is the proof of Lemma˜4.18, because the corresponding proof in [polynomial_strikes_back] crucially relies on a fact that does not hold for our problem. We will discuss it in detail below.
@@ 4.1 The (dual) polynomial method
We now consider a property of a sequence of integers as a function $F:D\subseteq[R]_{0}^{N}\to\{-1,1\}$. The symbol 0 will play a special role that will be exhibited later on. Unfortunately one cannot just take the polynomial of those integers. The standard approach (see [Aaronson02]) is to encode $s=(s_{1},\dots,s_{N})\in[R]_{0}^{N}$ into binary variables $x=(x_{i,j})_{i\in[N],j\in[R]_{0}}\in\{-1,1\}^{N(R+1)}$ encoding whether $s_{i}=j$ as follows: $x_{i,j}=-1$ if $s_{i}=j$, and $x_{i,j}=1$ otherwise. Let $H_{b}^{N(R+1)}\subseteq\{-1,1\}^{N(R+1)}$ be the set of all possible encodings of vectors $s$, that is, for every $i\in[N]$ there is exactly one $j\in[R]_{0}$ such that $x_{i,j}=-1$.
@@
But before doing this, we are going to relax the constraints on the domain $D_{b}$ in the case of symmetric functions, while we decrease its dimension. When $F$ is symmetric (i.e., $F(s)=F(s\circ\pi_{N})$ for any permutation $\pi_{N}$ of $[N]$), one can instead define a function $F_{\leq N}$ with weaker constraints by removing the variables corresponding to the symbol 0. Define $H_{\leq N}^{NR}$ as the set of length-$(NR)$ binary vectors with Hamming weight at most $N$. Given any $x\in H_{\leq N}^{NR}$, we define its frequency vector $z(x)=(z_{0},z_{1},\dots,z_{R})$ with $z_{j}=\#\{i:x_{ij}=-1\}$, for $1\leq j\leq R$, and $z_{0}=N-z_{1}-\ldots-z_{R}$. From the vector $z(x)$, one can define a valid sequence of integers $s(x)\in[R]_{0}^{N}$: it can be any sequence from $R]_{0}^{N}$ that has frequency vector $z(x)$. Now we can define $F_{\leq N}$ on domain $D_{\leq N}$ as
@@
Lemma 4.7 ([Aaronson02, ambainis_poly],[Bun_Thaler_20, Theorem 3.9]).
@@
Let $F:D\subseteq[R]_{0}^{N}\to\{-1,1\}$ be . Define $F_{\leq N}:D_{\leq N}\to\{-1,1\}$ as above. Then ${\mathrm{dpdeg}}_{\delta}(F_{\leq N})\geq\Delta$ iff there exists a function $\psi:\{-1,1\}^{NR}\to\mathbb{R}$ such that
@@ 4.2 Preparation
Technically, the problem we use in the proof of Theorem˜4.1 is slightly more restricted than $k$-collision-freeness: we want to distinguish no $k$-collision from many distinct collisions of size at least $k$.
@@
Let $\gamma\in(0,1)$. The function $\mathrm{Collision}^{k,\gamma}_{N,R}:D_{\mathrm{Collision}^{k,\gamma}_{N,R}}\subset[R]^{N}\to\{-1,1\}$ is defined by $\mathrm{Collision}^{k,\gamma}_{N,R}(s)=-1$ if no integer occurs at least $k$ times in $s$, $\mathrm{Collision}^{k,\gamma}_{N,R}(s)=1$ if there are more than $\gamma R$ distinct integers that occur at least $k$ times in $s$, and it is undefined otherwise.
@@
To prove a bound on the $\mathrm{Collision}$ function, we will actually relate it to the composition of two more elementary functions $g,h$, where by composition we mean $(g\circ h)(x)=g(h(x_{1}),\dots,h(x_{n}))$ (where $x=(x_{1},\dots,x_{n})$ and each $x_{i}$ is a binary vector of appropriate dimension). Let us define (i) the threshold function $\mathrm{THR}_{N}^{k}:\{-1,1\}^{N}\to\{-1,1\}$ which is 1-1 if the input bitstring contains at least $k$ many 1-1s, and it is 1 otherwise; and (2) the gap version of $\mathrm{OR}$, that is $\mathrm{GapOR}_{R}^{\gamma}:D_{\mathrm{GapOR}_{R}^{\gamma}}\subset\{-1,1\}^{R}\to\{-1,1\}$ which takes value 1 if the input is $^{R}, 1-1$ if the input contains at least $\gamma R$ many 1-1s, and is undefined otherwise. We show that the double-promise approximate degree of $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ lower bounds the quantum query complexity of the collision problem.
@@
Let $k\geq 3$, $0<\gamma<1$, $\delta>0$ and $c>2$ be constants such that $N/c\leq R\leq N/2$. If the double-promise $\delta$-approximate degree of $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ on domain further restricted to $H_{\leq N}^{NR}$ is at least $\Delta$, then every quantum algorithm computing $\mathrm{Collision}^{k,\gamma/c}_{N,N}$ with error $\delta/2$ must require at least $\Delta/2$ queries.
Before proving this lemma, we prove some helper propositions. In order to apply the dual polynomial method for partial functions, we start by proving that $\mathrm{Collision}^{k,\gamma^{\prime}}_{N,R^{\prime}}$ is at least as hard as a very similar problem. We introduce a “dummy-augmented” version $\mathrm{dCollision}_{N,R}^{k,\gamma}:D_{\mathrm{dCollision}_{N,R}^{k,\gamma}}\subseteq[R]_{0}^{N}\to\{-1,1\}$ of the problem $\mathrm{Collision}_{N,R}^{k,\gamma}$ for the purpose of proving Lemma˜4.10, where now the input sequence can have integer 0, but those 0s are just ignored when they occur. We show that it is enough to prove a lower bound for this second version.
@@
Let $k\geq 3$, $0<\gamma<1$ and $c>2$ be constants such that $N/c\leq R\leq N/2$. Then $\mathrm{dCollision}^{k,\gamma}_{N,R}$ can be reduced to $\mathrm{Collision}^{k,\gamma/c}_{N,N}$.
@@
An input to $\mathrm{dCollision}^{k,\gamma}_{N,R}$ is a sequence $s=(s_{1},\dots,s_{N})$ where each $s_{i}\in[R]_{0}$. Let us define a family of functions $T_{i}$ that map from $R]_{0}$ to $R^{\prime}]$ for $R^{\prime}=R+\lceil N/2\rceil$: $T_{i}(s)=s$ if $s>0$ and $T_{i}(0)=R+\lceil i/2\rceil$.
@@
We are now ready to give the proof of Lemma˜4.10.
Proof of Lemma˜4.10.
By Proposition˜4.11, instead of $\mathrm{Collision}^{k,\gamma/c}_{N,N}$ we can consider $\mathrm{dCollision}^{k,\gamma}_{N,R}$ (with the appropriate parameters) to show a lower bound. By Proposition˜4.12, we can use Lemma˜4.7 to relate the query complexity of $\mathrm{dCollision}^{k,\gamma}_{N,R}$ to the double-promise degree of $\mathrm{GapOR}_{R}^{\gamma}\circ\mathrm{THR}_{N}^{k}$ with domain further restricted to $H_{\leq N}^{NR}$. ∎